OpenStudy (captbrazil):

Factor the expression completely: (x-1)^7/2 - (x-1)^3/2

OpenStudy (evoker):

Start by pulling out the common factor

OpenStudy (captbrazil):

(u^2 - 1)^7/2 - (u^2 - 1)^3/2 = (u^7 - 1) - (u^6 - 1) where do I go from here?

OpenStudy (evoker):

I would pull out the common factor of (x-1)^3/2

OpenStudy (evoker):

This yields (x-1)^3/2 * ((x-1)^4-1)

OpenStudy (evoker):

Oh quick check is this \[(x-1)^{7/2}-(x-1)^{3/2}\]

OpenStudy (evoker):

or \[\frac{ (x-1)^7 }{ 2}-\frac{ (x-1)^3 }{ 2}\]

OpenStudy (captbrazil):

it is the first one

OpenStudy (evoker):

Ah ok my mistake in that case the common factor is (x-1)^(3/2)

OpenStudy (evoker):

so (x-1)^(3/2) * ((x-1)^(4/2)-1)

OpenStudy (evoker):

or (x-1)^(3/2)* ((x-1)^2-1)

OpenStudy (captbrazil):

hmm ok, that makes more sense

OpenStudy (captbrazil):

so are we all looking for is a common factor, dont have to replace with u^2?

OpenStudy (captbrazil):

x = u^2 that is

OpenStudy (evoker):

no not necessary

OpenStudy (evoker):

though probably need to expand that second term and see if we can factor

OpenStudy (evoker):

second term is x^2-2x so you can factor it into x(x-2)